### v1

@@ Abstract
Models like GPT-4o enable real-time interaction with large language models (LLMs) through speech, significantly enhancing user experience compared to traditional text-based interaction. However, there is still a lack of exploration on how to build speech interaction models based on open-source LLMs. To address this, we propose LLaMA-Omni, a novel model architecture designed for low-latency and high-quality speech interaction with LLMs. LLaMA-Omni integrates a pretrained speech encoder, a speech adaptor, an LLM, and a streaming speech decoder. It eliminates the need for speech transcription, and can simultaneously generate text and speech responses directly from speech instructions with extremely low latency. We build our model based on the latest Llama-3.1-8B-Instruct model. To align the model with speech interaction scenarios, we construct a dataset named InstructS2S-200K, which includes 200K speech instructions and corresponding speech responses. Experimental results show that compared to previous speech-language models, LLaMA-Omni provides better responses in both content and style, with a response latency as low as 226ms. Additionally, training LLaMA-Omni takes less than 3 days on just 4 GPUs, paving the way for the efficient development of speech-language models in the future. [^1] Code and model are available at https://github.com/ictnlp/LLaMA-Omni.
@@ 1 Introduction
In this paper, we propose a novel model architecture, LLaMA-Omni, which enables low-latency and high-quality interaction with LLMs. LLaMA-Omni consists of a speech encoder, a speech adaptor, an LLM, and a streaming speech decoder. The user’s speech instruction is encoded by the speech encoder followed by the speech adaptor, and then input into the LLM. The LLM decodes the text response directly from the speech instruction, without first transcribing the speech into text. The speech decoder is a non-autoregressive (NAR) streaming Transformer (Ma et al., 2023), which takes the output hidden states of the LLM as input and uses connectionist temporal classification (CTC; Graves et al., 2006a) to predict the sequence of discrete units corresponding to the speech response. During inference, as the LLM autoregressively generates the text response, the speech decoder simultaneously generates the corresponding discrete units. To better align with the characteristics of speech interaction scenarios, we construct a dataset named InstructS2S-200K by rewriting existing text instruction data and performing speech synthesis. Experimental results show that LLaMA-Omni can simultaneously generate high-quality text and speech responses with a latency as low as 226ms. Additionally, compared to previous speech-language models like SpeechGPT (Zhang et al., 2023), LLaMA-Omni significantly reduces the required training data and computational resources, enabling the efficient development of powerful speech interaction models based on the latest LLMs.
@@ 2.2 Speech Adaptor
To enable the LLM to comprehend the input speech, we incorporate a trainable speech adaptor $\mathcal{A}$ that maps the speech representations into the embedding space of the LLM. Following  Ma et al. (2024c), our speech adaptor first downsamples the speech representations $\mathbf{H}$ to reduce the sequence length. Specifically, every $k$ consecutive frames are concatenated along the feature dimension:
@@ 2.3 Large Language Model
We use Llama-3.1-8B-Instruct [^3] https://huggingface.co/meta-llama/Meta-Llama-3.1-8B-Instruct (Dubey et al., 2024) as the LLM $\mathcal{M}$, which is currently the state-of-the-art open-source LLM. It has strong reasoning capabilities and is well-aligned with human preferences. The prompt template $\mathcal{P}(\cdot)$ is shown in Figure 3. The speech representation sequence $\mathbf{S}$ is filled into the position corresponding to <speech>, and then the entire sequence $\mathcal{P}(\mathbf{S})$ is input into the LLM. Finally, the LLM autoregressively generates the text response $Y^{T}=[y^{T}_{1},...,y^{T}_{M}]$ directly based on the speech instruction and is trained using cross-entropy loss:
@@ 2.4 Speech Decoder
For the speech response $Y^{S}$, we first follow  Zhang et al. (2023) to discretize the speech into discrete units. Specifically, we use the pretrained HuBERT (Hsu et al., 2021) model to extract continuous representations of the speech, and then convert these representations into discrete cluster indices using a K-means model. Subsequently, consecutive identical indices are merged into a single unit, resulting in the final discrete unit sequence $Y^{U}=[y^{U}_{1},...,y^{U}_{L}],y^{U}_{i}\in\{0,1,...,K-1\},\forall 1\leq i\leq L$, where $K$ is the number of clusters, and $L$ is the length of discrete unit sequence. The discrete units can be converted into waveform with an additional unit-based vocoder $\mathcal{V}$ (Polyak et al., 2021).
@@
<|begin_of_text|><|start_header_id|>system<|end_header_id|>
You are a helpful language and speech assistant. You are able to
understand the speech content that the user provides, and assist the
user with a variety of tasks using natural language.<|eot_id|>
<|start_header_id|>user<|end_header_id|>
<speech>
Please answer the questions in the user’s input speech.<|eot_id|>
<|start_header_id|>assistant<|end_header_id|>
Figure 3: Prompt template of LLaMA-Omni.
@@ 2.6 Inference
During inference, the LLM autoregressively generates the text response based on the speech instruction. Meanwhile, since our speech decoder uses causal attention, once the LLM generates a text response prefix $Y^{T}_{\leq i}$, the corresponding upsampled hidden states $\widehat{\mathbf{C}}_{\leq\lambda\cdot i}$ can be fed into the speech decoder to generate a partial alignment $A_{\leq\lambda\cdot i}$, which in turn yields the discrete units corresponding to the generated text prefix. To further enable streaming synthesis of speech waveforms, when the number of generated units reaches a pre-defined chunk size $\Omega$, we input this unit segment into the vocoder to synthesize a speech segment, which is then immediately played to the user. As a result, users can start listening to the speech response without waiting for the complete text response to be generated, ensuring low response latency that is not affected by the length of the text response. Algorithm 1 describes the above process. Additionally, since the speech decoder uses non-autoregressive modeling, the alignment corresponding to each text token $y^{T}_{i}$, specifically $A_{\lambda\cdot(i-1)+1:\lambda\cdot i}$, is generated in parallel within the chunk. Therefore, the decoding speed for generating both text and speech simultaneously is not significantly different from the speed of generating text alone.
@@ 3 Construction of Speech Instruction Data: InstructS2S-200K
In speech interactions, existing responses from text instructions are not suitable for direct use as speech instruction responses. This is because, in text-based interactions, models tend to generate lengthy responses, using complex sentences and possibly including non-verbal elements like ordered lists or parentheses. However, in speech interactions, concise yet informative responses are typically preferred (Anonymous, 2024). Therefore, we use the Llama-3-70B-Instruct model to generate responses for speech instructions according to the following rules: (1) The response should not contain content that cannot be synthesized by the TTS model, such as parentheses, ordered lists, etc. (2) The response should be very concise and to the point, avoiding lengthy explanations. The prompt can be found in Appendix A.
@@
For the basic text instructions, we collect around 50K instructions from the Alpaca dataset [^7] https://huggingface.co/datasets/tatsu-lab/alpaca (Taori et al., 2023), which covers a wide range of topics. Additionally, we gather around 150K instructions from the UltraChat dataset [^8] https://github.com/thunlp/UltraChat (Ding et al., 2023), which primarily consist of questions about the world. Note that UltraChat is a large-scale multi-turn conversation dataset, but we only select the first 150K entries and use only the first-round instruction. Using the above datasets and data processing pipeline, we ultimately obtain 200K speech instruction data, referred to as InstructS2S-200K.
@@ 4.1 Experimental Setups
For the training data, we use the InstructS2S-200K dataset mentioned in Section 3, which includes 200K speech instruction data. To extract discrete units corresponding to the target speech, we use a pre-trained K-means quantizer [^9] https://dl.fbaipublicfiles.com/hubert/mhubert_base_vp_en_es_fr_it3_L11_km1000.bin, which has learned 1000 clusters from the HuBERT features. The pretrained HiFi-GAN vocoder (Kong et al., 2020; Polyak et al., 2021) is used to synthesize discrete units into waveform. For the evaluation data, we select two subsets from Alpaca-Eval [^10] https://github.com/tatsu-lab/alpaca_eval (Li et al., 2023): helpful_base and vicuna, as their questions are more suitable for speech interaction scenarios. We remove questions related to math and code, resulting in a total of 199 instructions. To obtain the speech version, we use the CosyVoice-300M-SFT model to synthesize the instructions into speech. We refer to this test set as InstructS2S-Eval in the following sections.
@@
We use the encoder of Whisper-large-v3 as the speech encoder, and use Llama-3.1-8B-Instruct as the LLM. The speech adapter performs a 5$\times$ downsampling on the speech representations. The speech decoder consists of 2 Transformer layers with the same architecture as LLaMA, with a hidden dimension of 4096, 32 attention heads, and a feed-forward network dimension of 11008, which contains 425M parameters. The upsample factor $\lambda$ is set to 25. For the minimum unit chunk size $\Omega$ input to the vocoder, we set $\Omega=+\infty$ in the main experiment, meaning we wait for the entire unit sequence to be generated before inputting it to the vocoder for speech synthesis. In subsequent experiments, we will analyze how adjusting the value of $\Omega$ can control response latency, as well as the trade-off between latency and speech quality.
@@ 4.2 Evaluation
Since LLaMA-Omni can generate both text and speech responses based on speech instructions, we evaluate the model’s performance on two tasks: speech-to-text instruction-following (S2TIF) and speech-to-speech instruction-following (S2SIF). We use greedy search to ensure reproducible experimental results. The model is evaluated from the following aspects:
@@
To evaluate the model’s ability to follow speech instructions, we use GPT-4o (OpenAI, 2024) to score the model’s responses. For the S2TIF task, scoring is based on the transcribed text of the speech instructions and the model’s text response. For the S2SIF task, we first transcribe the model’s speech responses into text using the Whisper-large-v3 model, and then score it in the same manner as the S2TIF task. GPT-4o gives scores on two aspects: content and style. The content score evaluates whether the model’s response adequately addresses the user’s instruction, while the style score assesses whether the model’s response style is suitable for speech interaction scenarios. The detailed prompt can be found in Appendix A.
Speech-Text Alignment
To evaluate the alignment between text responses and speech responses, we use the Whisper-large-v3 model to transcribe the speech responses into text, and then calculate the Word Error Rate (WER) and Character Error Rate (CER) between the transcribed text and the text response. We refer to these metrics as ASR-WER and ASR-CER, respectively.
Speech Quality
To evaluate the quality of the generated speech, we utilize a Mean Opinion Score (MOS) prediction model called UTMOS [^11] https://github.com/tarepan/SpeechMOS (Saeki et al., 2022), which is capable of predicting the MOS score of the speech to assess its naturalness. We refer to this metric as the UTMOS score.
Response Latency
The latency is a key metric for speech interaction models, referring to the time interval between the input of a speech instruction and the start of the speech response, which has a significant impact on user experience. Additionally, we calculate the number of words already generated in the text response when the speech response begins, referred to as the #lagging word.
@@ 4.3 Baseline Systems
SALMONN (+TTS)
SALMONN (Tang et al., 2024) is a LLM capable of accepting speech and audio inputs and responding with text, enabling it to perform the S2TIF task. For the S2SIF task, we add a VITS TTS model after SALMOON to generate speech responses in a cascaded manner.
Qwen2-Audio (+TTS)

### v2

@@ Abstract
Models like GPT-4o enable real-time interaction with large language models (LLMs) through speech, significantly enhancing user experience compared to traditional text-based interaction. However, there is still a lack of exploration on how to build speech interaction models based on open-source LLMs. To address this, we propose LLaMA-Omni, a novel end-to-end model architecture designed for low-latency and high-quality speech interaction with LLMs. LLaMA-Omni integrates a pretrained speech encoder, a speech adaptor, an LLM, and a streaming speech decoder. It eliminates the need for speech transcription, and can simultaneously generate text and speech responses directly from speech instructions with extremely low latency. We build our model based on the latest Llama-3.1-8B-Instruct model. To align the model with speech interaction scenarios, we construct a dataset named InstructS2S-200K, which includes 200K speech instructions and corresponding speech responses, with a style that better matches the characteristics of speech interaction scenarios. Experimental results show that compared to previous speech-language models, LLaMA-Omni provides better responses in both content and style, with a response latency as low as 236ms. Additionally, training LLaMA-Omni takes less than 3 days on just 4 GPUs, paving the way for the efficient development of speech-language models in the future. [^1] Code: https://github.com/ictnlp/LLaMA-Omni
      Model: https://huggingface.co/ICTNLP/Llama-3.1-8B-Omni
      Audio Samples: https://ictnlp.github.io/llama-omni-demo/
@@ 1 Introduction
In this paper, we propose a novel model architecture, LLaMA-Omni, which enables low-latency and high-quality interaction with LLMs. LLaMA-Omni consists of a speech encoder, a speech adaptor, an LLM, and a streaming speech decoder. The user’s speech instruction is encoded by the speech encoder followed by the speech adaptor, and then input into the LLM. The LLM decodes the text response directly from the speech instruction, without first transcribing the speech into text. The speech decoder is a non-autoregressive (NAR) streaming Transformer (Ma et al., 2023), which takes the output hidden states of the LLM as input and uses connectionist temporal classification (CTC; Graves et al., 2006a) to predict the sequence of discrete units corresponding to the speech response. During inference, as the LLM autoregressively generates the text response, the speech decoder simultaneously generates the corresponding discrete units. To better align with the characteristics of speech interaction scenarios, we construct a dataset named InstructS2S-200K by rewriting existing text instruction data and performing speech synthesis. Experimental results show that LLaMA-Omni can simultaneously generate high-quality text and speech responses with a latency as low as 236ms. Additionally, compared to previous speech-language models like SpeechGPT (Zhang et al., 2023), LLaMA-Omni significantly reduces the required training data and computational resources, enabling the efficient development of powerful speech interaction models based on the latest LLMs.
@@ 2.2 Speech Adaptor
To enable the LLM to comprehend the input speech, we incorporate a trainable speech adaptor $\mathcal{A}$ that maps the speech representations into the embedding space of the LLM. Following  Ma et al. (2024b), our speech adaptor first downsamples the speech representations $\mathbf{H}$ to reduce the sequence length. Specifically, every $k$ consecutive frames are concatenated along the feature dimension:
@@ 2.3 Large Language Model
We use Llama-3.1-8B-Instruct [^3] https://huggingface.co/meta-llama/Meta-Llama-3.1-8B-Instruct (Dubey et al., 2024) as the LLM $\mathcal{M}$, which is currently the state-of-the-art open-source LLM. It has strong reasoning capabilities and is well-aligned with human preferences. The prompt template $\mathcal{P}(\cdot)$ is shown in Appendix A. The speech representation sequence $\mathbf{S}$ is filled into the position corresponding to <speech>, and then the entire sequence $\mathcal{P}(\mathbf{S})$ is input into the LLM. Finally, the LLM autoregressively generates the text response $Y^{T}=[y^{T}_{1},...,y^{T}_{M}]$ directly based on the speech instruction and is trained using cross-entropy loss:
@@ 2.4 Speech Decoder
For the speech response $Y^{S}$, we first follow  Zhang et al. (2023) to discretize the speech into discrete units. Specifically, we use the pretrained HuBERT (Hsu et al., 2021) model to extract continuous representations of the speech, and then convert these representations into discrete cluster indices using a K-means model. Subsequently, consecutive identical indices are merged into a single unit, resulting in the final discrete unit sequence $Y^{U}=[y^{U}_{1},...,y^{U}_{L}],y^{U}_{i}\in\{0,1,...,K-1\},\forall 1\leq i\leq L$, where $K$ is the number of clusters, and $L$ is the length of discrete unit sequence. To synthesize waveforms based on discrete units, we adopt a unit-based HiFi-GAN vocoder with a duration predictor (Polyak et al., 2021). It first predicts the duration of each discrete unit and repeats them to model prosody, and then generates the waveform based on discrete units.
@@ 2.6 Inference
During inference, the LLM autoregressively generates the text response based on the speech instruction. Meanwhile, since our speech decoder uses causal attention, once the LLM generates a text response prefix $Y^{T}_{\leq i}$, the corresponding upsampled hidden states $\widehat{\mathbf{C}}_{\leq\lambda\cdot i}$ can be fed into the speech decoder to generate a partial alignment $A_{\leq\lambda\cdot i}$, which in turn yields the discrete units corresponding to the generated text prefix. To further enable streaming synthesis of speech waveforms, when the number of generated units reaches a pre-defined chunk size $\Omega$, we input this unit segment into the vocoder to synthesize a speech segment, which is then immediately played to the user. As a result, users can start listening to the speech response without waiting for the complete text response to be generated, ensuring low response latency that is not affected by the length of the text response. Algorithm 1 describes the above process. Additionally, since the speech decoder uses non-autoregressive modeling, the alignment corresponding to each text token $y^{T}_{i}$, specifically $A_{\lambda\cdot(i-1)+1:\lambda\cdot i}$, is generated in parallel within the chunk. Therefore, the decoding speed for generating both text and speech responses simultaneously is not significantly different from the speed of generating text response alone.
@@ 3 Construction of Speech Instruction Data: InstructS2S-200K
In speech interactions, existing responses from text instructions are not suitable for direct use as speech instruction responses. This is because, in text-based interactions, models tend to generate lengthy responses, using complex sentences and possibly including non-verbal elements like ordered lists or parentheses. However, in speech interactions, concise yet informative responses are typically preferred (Cho et al., 2024). Therefore, we use the Llama-3-70B-Instruct model to generate responses for speech instructions according to the following rules: (1) The response should not contain content that cannot be synthesized by the TTS model, such as parentheses, ordered lists, etc. (2) The response should be very concise and to the point, avoiding lengthy explanations. The prompt can be found in Appendix A.
@@
Table 1: Statistical information of the InstructS2S-200K dataset.
Statistic Value
Speech Instruction Duration 418h
Speech Response Duration 1058h
Avg. Speech Instruction Duration 7.5s
Avg. Speech Response Duration 19.0s
Avg. Text Instruction Length 21.7
Avg. Text Response Length 39.5
Avg. Unit Sequence Length 553.6
For the basic text instructions, we collect around 50K instructions from the Alpaca dataset [^7] https://huggingface.co/datasets/tatsu-lab/alpaca (Taori et al., 2023), which covers a wide range of topics. Additionally, we gather around 150K instructions from the UltraChat dataset [^8] https://github.com/thunlp/UltraChat (Ding et al., 2023), which primarily consist of questions about the world. Note that UltraChat is a large-scale multi-turn conversation dataset, but we only select the first 150K entries and use only the first-round instruction. Using the above datasets and data processing pipeline, we ultimately obtain 200K speech instruction data, referred to as InstructS2S-200K. The detailed statistical information is listed in Table 1.
@@ 4.1 Experimental Setups
For the training data, we use the InstructS2S-200K dataset mentioned in Section 3, which includes 200K speech instruction data. To extract discrete units corresponding to the target speech, we use a pre-trained K-means quantizer [^9] https://dl.fbaipublicfiles.com/hubert/mhubert_base_vp_en_es_fr_it3_L11_km1000.bin, which has learned 1000 clusters from the HuBERT features. The pretrained HiFi-GAN vocoder [^10] https://dl.fbaipublicfiles.com/fairseq/speech_to_speech/vocoder/code_hifigan/mhubert_vp_en_es_fr_it3_400k_layer11_km1000_lj/g_00500000 (Kong et al., 2020; Polyak et al., 2021) is used to synthesize discrete units into waveform. For the evaluation data, we select two subsets from Alpaca-Eval [^11] https://github.com/tatsu-lab/alpaca_eval (Li et al., 2023): helpful_base and vicuna, as their questions are more suitable for speech interaction scenarios. We remove questions related to math and code, resulting in a total of 199 instructions. To obtain the speech version, we use the CosyVoice-300M-SFT model to synthesize the instructions into speech. We refer to this test set as InstructS2S-Eval in the following sections.
@@
We use the encoder of Whisper-large-v3 as the speech encoder, and use Llama-3.1-8B-Instruct as the LLM. The speech adapter performs a 5$\times$ downsampling on the speech representations. The speech decoder consists of 2 Transformer layers with the same architecture as LLaMA, with a hidden dimension of 4096, 32 attention heads, and a feed-forward network dimension of 11008, which contains 425M parameters. The upsample factor $\lambda$ is set to 25. For the minimum unit chunk size $\Omega$ input to the vocoder, we set $\Omega=+\infty$ in the offline scenario, meaning we wait for the entire unit sequence to be generated before inputting it to the vocoder for speech synthesis. In the streaming scenario, we adjust the value of $\Omega$ within the range of [10, 20, 40, 60, 80, 100] to control the response latency of the model.
@@ 4.2 Evaluation
Since LLaMA-Omni can generate both text and speech responses based on speech instructions, we evaluate the model’s performance on two tasks: speech-to-text instruction-following (S2TIF) and speech-to-speech instruction-following (S2SIF). We use greedy search to ensure reproducible experimental results. For the S2SIF task, we further evaluate the model in two scenarios: offline and streaming. In the offline scenario, the model generates the text response first, and then synthesizes the complete speech. In the streaming scenario, the speech response is generated simultaneously with the text response. We use the following metrics to evaluate the model:
@@
To evaluate the model’s ability to follow speech instructions, we use GPT-4o (OpenAI, 2024) to score the model’s responses. For the S2TIF task, scoring is based on the transcribed text of the speech instructions and the model’s text response. For the S2SIF task, we first transcribe the model’s speech responses into text using the Whisper-large-v3 model, and then score it in the same manner as the S2TIF task. GPT-4o gives a score between 1 and 5 based on factors such as helpfulness, relevance, fluency, and suitability for speech interaction. The detailed prompt for evaluation can be found in Appendix A.
ASR-WER
To evaluate the alignment between text and speech responses, we use the Whisper-large-v3 model to transcribe the speech responses into text, and then calculate the Word Error Rate (WER) between the transcribed text and the text response, which is referred to as ASR-WER.
UTMOS
To evaluate the quality of the generated speech, we utilize a Mean Opinion Score (MOS) prediction model called UTMOS [^12] https://github.com/tarepan/SpeechMOS (Saeki et al., 2022), which is capable of predicting the MOS score of the speech to assess its naturalness.
Latency
The response latency is a key metric for speech interaction models, referring to the time interval between the input of a speech instruction and the start of the speech response, which has a significant impact on user experience. We measure the latency on 1 NVIDIA L40 GPU.
Speech Rate
To measure the speech rate of the generated speech, we use the metric Words Per Second (WPS), which represents the average number of words per second in the generated speech.
@@ 4.3 Baseline Systems
SALMONN + Orca
SALMONN (Tang et al., 2024) is a LLM capable of accepting speech and audio inputs and responding with text, enabling it to perform the S2TIF task. For the S2SIF task, we add a TTS model called Orca [^13] https://github.com/Picovoice/orca after SALMONN. Orca is an industrial TTS model that supports both streaming and offline speech synthesis, delivering excellent performance. This integrated system enables speech synthesis to begin concurrently with the generation of the text response.
Qwen2-Audio + Orca